# Stein Random Feature Regression

### v1

@@ 4.4 Mixture Stein Random Feature Regression
2:     Compute gradient $\nabla_{{\bm{\mathbf{\Omega}}}_{m}}\log p(D|{\bm{\mathbf{\Omega}}}_{m})p({\bm{\mathbf{\Omega}}}_{m}^{t})$ with $p(D|{\bm{\mathbf{\Omega}}}_{m})$ from (9) for all $m\in M$.
3:     for each frequency matrix ${\bm{\mathbf{\Omega}}}_{m},m\in M$ do
4:        for each frequency matrix ${\bm{\mathbf{\Omega}}}_{j},j\in M$ do
@@
which represents the gradient of the SSGP likelihood (9) given the RFF frequency matrix ${\bm{\mathbf{\Omega}}}_{m}$ and the sum of a frequency prior $p({\bm{\omega}})$ over the component frequencies $\{{\bm{\omega}}_{i,m}\}_{i=1}^{R}={\bm{\mathbf{\Omega}}}_{m}$. We define the inter-particle kernel as a matrix-valued kernel between the row vectors of each matrix:
@@
M-SRFR is summarized in Algorithm 1, and we provide further implementation and gradient calculation details in Appendix B.4. The construction of ${\bm{\mathbf{\Omega}}}_{m}$ via frequencies $\{{\bm{\omega}}_{i,m}\}_{i=1}^{R}$ implies that gradient updates to ${\bm{\mathbf{\Omega}}}_{m}$ directly modify its constituent frequencies. Collectively, the parameters of the M-SRFR model form a tensor with dimensions $M\times R\times d$.
For making predictions with M-SRFR, summarized in Algorithm 2, we combine the $M$ individual mixture predictive means and covariances into a single predictive distribution. We do so using properties of Gaussian mixtures, which allows for calculation of an overall mean and covariance as a uniformly-weighted aggregate of the mixture components. An example comparison of our method and traditional RFF learning is presented in Figure 1.
@@
1:  for each frequency matrix ${\bm{\mathbf{\Omega}}}_{m},m\in M$ do
@@ 5.1 Stein Random Features for Kernel Approximation
In a comparative experiment focused on the Gaussian (RBF) kernel, we evaluate the approximation quality of randomized kernel Gram matrices $\mathbf{K}$ using RFFs with varied sampling techniques as well as other common low-rank kernel approximation methods. Specifically, we benchmark Matrix-SVGD [Wang et al., 2019] sampling against MC, QMC, orthogonal random features (ORF) [Yu et al., 2016], and Nystroem approximation (NYS) [Yang et al., 2012]. Results are demonstrated in Figure 2, where we use as a metric the Frobenius norm $\frac{||\mathbf{K}-\hat{\mathbf{K}}||}{||\mathbf{K}||}$ between Gram approximation $\hat{\mathbf{K}}$ and true Gram matrix $\mathbf{K}$.
The results underscore SVGD’s strong approximation capabilities over other sampling techniques across different ranks $R$, and notably, SVGD outperforms data-dependent methods like Nystroem as $R$ increases. Across data dimensionality $d$, SVGD scales better than existing sampling based approaches, though not as efficiently as Nystroem. Nonetheless, given the challenges kernel methods face in high-dimensional spaces, practitioners likely resort to dimensionality reduction before applying kernel techniques in such high-dimensional settings.
@@ 5.2 UCI Regression Benchmarks
We evaluate M-SRFR on a variety of regression problems from the UCI data repository [Dua and Graff, 2017], with baselines of sparse variational Gaussian processes [Hensman et al., 2015] (SVGP), SSGPs with an RFF Gaussian kernel (SSGP-RBF), an SSGP with frequencies as hyperparameters, and an SSGP trained using the entropy-regularized functional kernel learning approach described in Section 2.4 (SSGP-SVGD), for which more details can be found in Wang and Liu [2019]. Additionally we introduce SSGP-$R^{*}$, which has $R^{*}=\sqrt[3]{MR^{3}}$ frequency samples, matching M-SRFR’s computational complexity for $Q$ mixture components.
All models and baselines are given a proper hyperparameter optimization treatment. The results in Table 1, where $Q$ represents the number of mixture components used in the M-SRFR model, highlight M-SRFR’s superior RMSE performance, particularly when contrasted with SSGP-$R^{*}$, underscoring the mixture approach’s advantage over simply enhancing RFF feature count. We employed wide Gaussian priors on the M-SRFR frequencies, and posit that specialized priors tailored to data characteristics may further enhance performance.
@@
NLPD, which measures uncertainty calibration, results vary. All models exhibit high variances across seeds, hence we do not bold results. M-SRFR is competitive on many datasets in NLPD but is outperformed by simpler baselines that tend to make wider and less mean-accurate predictions, as we demonstrate in Appendix D. The drop-off in performance from RMSE to NLPD for M-SRFR may be due to the non-standard predictive methodology defined in Algorithm 2, suggesting potential refinements for future investigations.
@@ 5.3 Large-Scale Ocean Modeling
Lastly, we evaluate our methods on a larger real-world problem using public data sourced from the AUSWAVE physics model produced by the Australian Bureau of Meteorology [2016]. The task is to predict significant wave height across the spatial domain, shown in Figure 3, using $N=5000$ randomly sampled locations with an input dimension $d=8$ model covariates and spatial coordinates.
We chose this setting as it is an inherently non-stationary domain, both due to the complexity of oceanographic modeling, as well as the fact that the distribution is not supported over the entire spatial domain. As such, we demonstrate the flexibility of M-SRFR to adapt to alternative kernel learning methodologies by introducing an M-SRFR variant that can handle non-stationarity through the use of deep kernels [Wilson et al., 2016]. Specifically, we jointly train a neural network with 3 hidden layers and 32 activations per layer to first project the data before an M-SRFR kernel mixture of RFFs is applied. The neural network is trained jointly through gradient descent with the M-SRFR RFF parameters, but does not receive the same “mixture” treatment – ie. all $M$ M-SRFR RFF kernels share the same network. We additionally an include an SSGP with a deep kernel, SSGP-Deep, as an additional baseline in order to demonstrate the unique benefit performing inference over kernel spectral measures.
The results, shown in Figure 3 (right), show that M-SRFR’s flexibility offers empirical benefit. Most the of the stationary baselines, including traditional M-SRFR, have difficulty adjusting to the non-stationary domain. However, the deep M-SRFR variant significantly outperforms even the deep SSGP variant, as well as all baselines, demonstrating that there is unique value to the mixture approach. When paired with the flexibility for integration with other kernel architectures such as the deep kernel, M-SRFR can adapt and perform well on a variety of problem domains with little change of methodology.

### v2

@@ 4.4 Mixture Stein Random Feature Regression
2:     Compute gradient $\nabla_{{\bm{\mathbf{\Omega}}}_{m}}\log p(D|{\bm{\mathbf{\Omega}}}_{m})p({\bm{\mathbf{\Omega}}}_{m})$ with $p(D|{\bm{\mathbf{\Omega}}}_{m})$ from (9) for all $m\in\{1,\dots,M\}$.
3:     for each ${\bm{\mathbf{\Omega}}}_{m},m\in\{1,\dots,M\}$ do
4:        for each ${\bm{\mathbf{\Omega}}}_{j},j\in\{1,\dots,M\}$ do
@@
which represents the gradient of the SSGP likelihood (9) given the RFF frequency matrix ${\bm{\mathbf{\Omega}}}_{m}$ and the sum of a frequency prior $p({\bm{\omega}})$ over the component frequencies $\{{\bm{\omega}}_{i,m}\}_{i=1}^{R}={\bm{\mathbf{\Omega}}}_{m}$. The inter-particle kernel is given by the kernel between the rows of each matrix of frequencies:
@@
M-SRFR is summarized in Algorithm 1, and we provide further implementation and gradient calculation details in Appendix B.4. The construction of ${\bm{\mathbf{\Omega}}}_{m}$ via frequencies $\{{\bm{\omega}}_{i,m}\}_{i=1}^{R}$ implies that gradient updates to ${\bm{\mathbf{\Omega}}}_{m}$ directly modify its constituent frequencies. Collectively, the parameters of the M-SRFR model form a tensor with dimensions $M\times R\times d$. An example comparison of M-SRFR and traditional RFF learning is presented in Figure 1.
M-SRFR performs approximate Bayesian posterior inference over $P(\pi)$ when the temperature parameter $\alpha=1$. However, when working with a small number of mixture components $M$, we observed empirical benefit to including $\alpha$ in the hyperparameter training routine to regulate the strength of the repulsive force.
For making predictions with M-SRFR, summarized in Algorithm 2, we combine the $M$ individual mixture predictive means and covariances into a single predictive distribution. We do so using properties of Gaussian mixtures, which allows for calculation of an overall mean and covariance as a uniformly-weighted aggregate of the mixture components.
@@
1:  for each frequency matrix ${\bm{\mathbf{\Omega}}}_{m},m\in\{1,\dots,M\}$ do
@@ 5.1 Stein Random Features for Kernel Approximation
In a comparative experiment focused on the Gaussian (RBF) kernel, we evaluate the approximation quality of randomized kernel Gram matrices $\mathbf{K}$ using RFFs with varied sampling techniques as well as other common low-rank kernel approximation methods. Specifically, we benchmark Matrix-SVGD [Wang et al., 2019] sampling against MC, QMC, orthogonal random features (ORF) [Yu et al., 2016], and Nyström approximation (NYS) [Yang et al., 2012]. Results are demonstrated in Figure 2, where we use as a metric the Frobenius norm $\frac{||\mathbf{K}-\hat{\mathbf{K}}||}{||\mathbf{K}||}$ between Gram approximation $\hat{\mathbf{K}}$ and true Gram matrix $\mathbf{K}$.
The results underscore SVGD’s strong approximation capabilities over other sampling techniques across different ranks $R$, and notably, SVGD outperforms data-dependent methods like Nystroem as $R$ increases. Across data dimensionality $d$, SVGD scales better than existing sampling based approaches, though not as efficiently as Nyström. Nonetheless, given the challenges kernel methods face in high-dimensional spaces, practitioners likely resort to dimensionality reduction before applying kernel techniques in such high-dimensional settings.
@@ 5.2 UCI Regression Benchmarks
We evaluate M-SRFR on a variety of regression problems from the UCI data repository [Dua and Graff, 2017], with baselines of sparse variational Gaussian processes [Hensman et al., 2015] (SVGP), SSGPs with an RFF Gaussian kernel (SSGP-RBF), an SSGP with frequencies as hyperparameters, and an SSGP trained using the entropy-regularized functional kernel learning approach described in Section 2.4 (SSGP-SVGD), for which more details can be found in Wang and Liu [2019]. Additionally we introduce SSGP-$R^{*}$, which has $R^{*}=\sqrt[3]{MR^{3}}$ frequency samples, matching M-SRFR’s computational complexity for $M$ mixture components.
All models and baselines are given a proper hyperparameter optimization treatment. The results in Table 1, where $M$ represents the number of mixture components used in the M-SRFR model, highlight M-SRFR’s superior RMSE performance, particularly when contrasted with SSGP-$R^{*}$, underscoring the mixture approach’s advantage over simply enhancing RFF feature count. We employed wide Gaussian priors on the M-SRFR frequencies, and posit that specialized priors tailored to data characteristics may further enhance performance.
@@
NLPD, which measures uncertainty calibration, results vary. All models exhibit high variance across seeds, hence we do not bold results. M-SRFR is competitive in NLPD on many datasets but is outperformed by simpler baselines that tend to make wider and less mean-accurate predictions, as we demonstrate in Appendix D. The drop-off in performance from RMSE to NLPD for M-SRFR may be due to the non-standard predictive methodology defined in Algorithm 2, suggesting potential refinements for future investigations.
@@ 5.3 Large-Scale Ocean Modeling
Lastly, we evaluate our methods on a real-world problem using public data sourced from the AUSWAVE physics model produced by the Australian Bureau of Meteorology [2016]. The task is to predict significant wave height across the spatial domain, shown in Figure 3, using $N=5000$ randomly sampled locations and an input dimension $d=8$ consisting of the spatial coordinates with additional physical model covariates.
We chose this setting as it is an inherently non-stationary domain, both due to the complexity of oceanographic modeling, as well as the fact that the distribution is not supported over the entire spatial domain. As such, we demonstrate the flexibility of M-SRFR to adapt to alternative kernel learning methodologies by introducing a non-stationary M-SRFR variant through the use of deep kernels [Wilson et al., 2016]. Specifically, we jointly train a neural network with 3 hidden layers and 32 activations per layer to first project the data before an M-SRFR kernel mixture of RFFs is applied. The neural network is trained jointly with the M-SRFR RFF parameters, but does not receive the same “mixture” treatment – ie. all $M$ M-SRFR kernels share the same input network.
In this sense, we are measuring M-SRFR’s ability to slot in as a modular component to alternative kernel learning schemes, and whether the benefits of the mixture approach extend to such a setting. We include an SSGP with a deep kernel, SSGP-Deep, as an additional baseline in order to differentiate the effect of the M-SRFR mixture from the neural network projection.
The results, shown in Figure 3 (right), show that M-SRFR’s flexibility offers significant benefit. Most the of the stationary baselines, including traditional M-SRFR, have difficulty adjusting to the non-stationary domain. However, the deep M-SRFR variant significantly outperforms even the SSGP-Deep variant, as well as all baselines, demonstrating that there is unique value to the mixture approach. These results highlight that with little change of methodology, M-SRFR can be injected into alternative kernel learning schemes to improve performance and flexibility.